# An operational approach to classifying measurement incompatibility

### v1

@@ II.1 Coarse-graining of outcomes
A set of measurements $\{M_{\overline{z}_{x}|x}\}$ is fully incompatible with respect to (w.r.t) coarse-graining if they remain incompatible after all possible nontrivial coarse-graining. That is, if the resultant set of measurements given by
$M_{z_{x}|x}=\sum_{\overline{z}_{x}}c_{x}(z_{x}|\overline{z}_{x})M_{\overline{z}_{x}|x},$ (3)
after all possible sets of nontrivial coarse-graining $\{c_{x}(z_{x}|\overline{z}_{x})\}$ pertaining to every setting $x$ is incompatible, then we call them fully incompatible.
Definition 2 ($d$-incompatible measurements w.r.t coarse-graining).
@@
A set of measurements $\{M_{\overline{z}_{x}|x}\}$ is $d$-incompatible w.r.t coarse-graining if they remain incompatible after all possible nontrivial coarse-graining that give rise to at least $d$ outcome measurements. In other words, if the resultant set of measurements
$M_{z_{x}|x}=\sum_{\overline{z}_{x}}c_{x}(z_{x}|\overline{z}_{x})M_{\overline{z}_{x}|x},\ z_{x}\in[k_{x}],k_{x}\geqslant d,\forall x$ (4)
after all possible coarse-graining resulting in at least $d$ outcome measurements is incompatible, then we call them $d$-incompatible.
@@
For example, the four outcome rank-one projective measurement pair, defined by the following (un-normalized) vectors
$M=\Big{\{}|0\rangle,|1\rangle,|2\rangle,|3\rangle\Big{\}}$ (5)
@@
$N=\Big{\{}|0+1\rangle,|0-1\rangle,|2+3\rangle,|2-3\rangle\Big{\}}$ (6)
is 3-incompatible, where the notation $|0+1\rangle=|0\rangle+|1\rangle,$ and similarly for others.
@@
A set of fully incompatible measurements is equivalent to 2-incompatible measurements w.r.t coarse-graining.
If a set of measurements is 2-incompatible, then it implies that the set remains incompatible after all possible coarse-graining of the outcomes such that the number of outcomes of each measurement in the newly formed set of measurements is greater than or equal to two. Also, the lowest number of outcomes of measurement is two for a nontrivial coarse-graining. Furthermore, if a set of measurements is $d$-incompatible, then, by definition, it is $n$-incompatible as well, where $n>d$, but the reverse is not true. This proves Observation 1.
@@
[…]
@@
The following condition is necessary but not sufficient for two rank-one projective measurements, defined by $\{|\psi_{i}\rangle\}$ and $\{|\phi_{j}\rangle\}$, to be fully incompatible w.r.t coarse-graining:
@@
The above condition is necessary and sufficient for two 3-dimensional rank-one projective measurements.
@@
First, note that for sharp measurement compatibility and commutativity are equivalent Heinosaari and Wolf (2010). Suppose $\exists i,j,$ such that $\langle\psi_{i}|\phi_{j}\rangle=0$. Consider coarse-graining of all other outcomes except $i$ and $j$ for the two measurements. Then, the resulting measurements will be $\{|\psi_{i}\rangle\!\langle\psi_{i}|,\mathbbm{1}-|\psi_{i}\rangle\!\langle\psi_{i}|\}$ and $\{|\phi_{j}\rangle\!\langle\phi_{j}|,\mathbbm{1}-|\phi_{j}\rangle\!\langle\phi_{j}|\}$, which commute with each other (i.e., the resulting measurements are compatible) since $\langle\psi_{i}|\phi_{j}\rangle=0$. Thus, if the measurements are fully incompatible, condition (8) holds.
@@
II.2 Convex mixing of measurements
The concept of convex mixing of measurements may be best understood by considering an example: suppose we have a measurement device that produces three different measurements $(M_{0},M_{1}\,\,\text{and}\,\,M_{2})$ and has two input switches. If the first switch is ON it performs measurement $M_{0}$ with probability $q$ and $M_{1}$ with probability $(1-q)$ and for the second switch, it always performs $M_{2}.$ Thus, in principle, the device produces two measurements: $M_{2}$ and a convex mixture of $M_{0}$ and $M_{1}.$
Suppose $M_{0}=\{M_{z|0}\},M_{1}=\{M_{z|1}\},M_{2}=\{M_{z|2}\}$ such that they have same number of outcomes $z\in[d]$. Now the measurement that is realized by a convex combination of $M_{0}$ and $M_{1}$ with some weightage $q\,\,\text{and}\,\,1-q$, where $q\in[0,1]$is defined as
$M_{(0,1,\pi)}=\{qM_{z|0}+(1-q)M_{\pi(z)|1}\},$ (11)
where $\pi(z)$ is any permutation on outcome $z$. Here $\pi(z)$ is a bijective function from the set $z\in[d]$ to itself. Even if $M_{2}$ is incompatible with $M_{0}$ and $M_{1}$ separately, $M_{2}$ is not necessarily incompatible with $M_{(0,1,\pi)}$ for all values of $q$.
@@ II.2 Convex mixing of measurements
[…]
Consider three unbiased qubit measurements,
$M_{i}:=\frac{1}{2}\left(\mathbbm{1}\pm\vec{n_{i}}\cdot\vec{\sigma}\right),\quad i=0,1,2,$ (12)
and $||\vec{n_{i}}||\leqslant 1$. These three measurements are fully incompatible w.r.t convex mixtures if and only if (see eq.(7) of Gühne et al. (2023)),
$||\vec{n}_{i}+q\vec{n}_{j}\pm(1-q)\vec{n}_{k}||+||\vec{n}_{i}-q\vec{n}_{j}\mp(1-q)\vec{n}_{k}||>2,$ (13)
@@
Observation 3.
If three qubit measurements (12) are such that $\vec{n}_{i}$ are in the same plane of the Bloch sphere, then they are not fully incompatible w.r.t convex mixtures.
@@
Note that permutation of the outcomes of $M_{i}$ in (12) is nothing but taking $-\vec{n}_{i}$ instead of $\vec{n}_{i}$. If the three $\vec{n}_{i}$ are in the same plane, then there exists at least one triple $(i,j,k)$ and permutations of outcomes such that
@@
for some $c\in(0,1],q\in[0,1]$. In other words, there exists a triple $(i,j,k)$ and permutation of outcomes such that $\vec{n}_{i}$ is expressed as a linear combination of $\vec{n}_{j}$ and $\vec{n}_{k}$ with non-negative coefficients (here $q/c$ and $(1-q)/c$), such that sum of those two non-negative coefficients is greater than or equal to 1. Substituting this into left hand side of (13) taking $+$ sign, we find
@@
Observation 4.
Suppose three qubit measurements (12) are such that $\vec{n}_{0}=\nu_{0}\hat{x}$, $\vec{n}_{1}=\nu_{1}\hat{y}$, $\vec{n}_{2}=\nu_{2}\hat{z}$ with $0\leq\nu_{1},\nu_{2},\nu_{3}\leq 1$, that is, they are the noisy version of Pauli observables,
$M_{i}:=\frac{1}{2}\left(\mathbbm{1}\pm\nu_{i}\sigma_{i}\right)=\nu_{i}\left(\frac{\mathbbm{1}\pm\sigma_{i}}{2}\right)+(1-\nu_{i})\frac{\mathbbm{1}}{2},$ (16)
taking $i=0,1,2$ and $(\sigma_{0},\sigma_{1},\sigma_{2})=(\sigma_{x},\sigma_{y},\sigma_{z})$. Then they are fully incompatible w.r.t convex mixture if and only if
@@
Thus, the r.h.s of (13) is greater than 2 for all values of $q$ if and only if
@@
Clearly, the three Pauli observables are fully incompatible w.r.t. convex mixture, and moreover, if we take an equal amount of noise $\nu_{0}=\nu_{1}=\nu_{2}=\nu$, then (17) implies $\nu>\sqrt{2/3}$.
We can generalize this notion for a set of $n$ measurements.
Definition 4 ($k$-incompatible measurements w.r.t. convex mixtures).
A set of measurements is $k$-incompatible w.r.t convex mixtures if after taking all possible convex mixtures over all possible $k$ partitions the resulting $k$ number of measurements from the set of $n$ measurements (with relabelling of outcomes) is incompatible.
Definition 5 (Fully incompatible measurements w.r.t. convex mixtures).
A set of measurements is fully incompatible w.r.t convex mixtures if it remains incompatible after taking all possible convex mixtures over all possible partitions of $n$ measurements (with relabelling of outcomes).
Theorem 2.
@@
Fully incompatible measurements w.r.t convex mixtures imply that every pair of measurements from that set is incompatible. The reverse implication does not hold.
@@
Consider a set of $n$ measurements, $M=\{M_{1},M_{2},M_{3},\ldots,M_{n}\}$, in which there is a pair of measurements $M_{i}$ and $M_{j}$ which are compatible. Now if we make two partitions where $M_{i}$ and $M_{j}$ are in different partitions, the convex mixture of measurements must be compatible. This is true for any compatible pair $M_{i}$ and $M_{j}$. Thus, if the measurements are fully incompatible w.r.t convex mixture, every pair of measurements must necessarily be incompatible.
The reverse is not true. Consider the three noisy Pauli measurements of Eq.(16). It can be shown by using semi-definite programming that if $0.71<\nu\leq 0.81$, the measurements are pairwise incompatible but it is not fully incompatible w.r.t convex mixture.
III Robustness under noise for different levels of incompatibility
@@
In this section, we analyze the role of noise on quantum measurements and study how the incompatibility properties depend on it. Due to the ubiquitous nature of noise, it is pertinent to study the extent to which noise could be tolerated by a set of measurements while still retaining their incompatibility. We take a noisy version of mutually unbiased bases measurements, $M_{i|x}\equiv\{\nu|\phi_{i}^{x}\rangle\langle\phi_{i}^{x}|+\frac{(1-\nu)}{d}\mathbbm{1}_{d}\},\,\,\text{where}\,\,\{|\phi_{i}^{x}\rangle\}$ form mutually unbiased bases measurements in $\mathbb{C}^{d}$. Here $\nu$ is the robustness parameter (or visibility parameter) and $(1-\nu)$ is the noise parameter, $0\leq\nu\leq 1$. When the noise parameter is zero (i.e., robustness, $\nu=1$), the measurements are fully incompatible, and when the noise parameter is one (i.e., the robustness, $\nu=0$), the measurements are trivial and compatible. Our aim is to obtain the critical value of the robustness parameter above which the measurements remain incompatible.
@@
To check the compatibility, i.e., the existence of a parent POVM, we use the method described in Gühne et al. (2023). This can be casted as a semi-definite programming (SDP) problem that takes a set of measurements $\{M_{z_{x}|x}\}$ and deterministic classical post-processings $p(z_{x}|x,\lambda)$ as input, and checks whether the measurements are compatible or not, subject to the constraints
@@
where $\mu$ is the optimization parameter. This method finds the maximum value of $\mu$ for each $\{p(z_{x}|x,\lambda)\}$. If this optimization returns a negative value of $\mu$, then the constraint of Eq.(22) can’t be fulfilled, which implies that the measurements $\{M_{z_{x}|x,\lambda}\}$ are incompatible. Otherwise, they are compatible.
III.1 Coarse-graining of outcomes
Dimension $3$. Let $\{P_{i}\}$ and $\{Q_{j}\},i,j\in[3]$ be two measurements acting on $\mathbb{C}^{3}$, where
$P_{i}=\nu|i\rangle\langle i|+(1-\nu)\frac{\mathbbm{1}}{3},$
$Q_{j}=\nu|\psi_{j}\rangle\langle\psi_{j}|+(1-\nu)\frac{\mathbbm{1}}{3},$ (23)

### v2

@@ III.1 Coarse-graining of outcomes
A set of measurements $\{M_{z|x}\}$ is fully incompatible w.r.t. coarse-graining if they remain incompatible after all possible nontrivial coarse-graining. That is if the resultant set of measurements given by Eq. (7) after all possible sets of nontrivial coarse-graining is incompatible, then we call them fully incompatible. Note that, the coarse-graining can be different for different settings $x$.
Definition 2 ($k$-incompatible measurements w.r.t. coarse-graining).
@@
A set of measurements $\{M_{z|x}\}$ is $k$-incompatible w.r.t. coarse-graining if they remain incompatible after all possible nontrivial coarse-graining that gives rise to at least $k$ outcome measurements. In other words, if the resultant set of measurements given by Eq. (7), where $\tilde{z}_{x}\in[d_{x}]$ and $d_{x}\geqslant k$ for all $x,$ is incompatible, then we call them $k$-incompatible.
@@
For example, the three outcome rank-one projective measurement pair, defined by the following (un-normalized) vectors
$M=\big{\{}|0\rangle,|1\rangle,|2\rangle\big{\}}$ (8)
@@
$N=\big{\{}|0\rangle+|1\rangle,|0\rangle-|1\rangle,|2\rangle\big{\}}$ (9)
is 3-incompatible, but not 2-incompatible since coarse-graining of the first two outcomes of these measurements yields compatible measurements.
@@
A set of fully incompatible measurements is equivalent to 2-incompatible measurements w.r.t. coarse-graining.
Proof.
[…]
@@
Consider two projective measurements, defined by $\{P_{i}\},\,\,\text{and}\,\,\{Q_{j}\}$, where $i\in[d]$ and $j\in[d^{\prime}]$. Let $\{\mathcal{M}_{k}\}_{k}$ be the set of all proper subsets of $[d]$, and $\{\mathcal{N}_{l}\}_{l}$ be the set of all proper subsets of $[d^{\prime}]$. Then these two measurements are fully incompatible w.r.t. coarse-graining if and only if
@@
The following condition is necessary but not sufficient for two rank-one projective measurements of dimension $\geq 4$, defined by $\{|\psi_{i}\rangle\}$ and $\{|\phi_{j}\rangle\}$, to be fully incompatible w.r.t. coarse-graining:
@@
However, the above condition is necessary and sufficient for two 3-dimensional rank-one projective measurements.
@@
First, note that for sharp measurement, compatibility and commutativity are equivalent Heinosaari and Wolf (2010). Suppose $\exists i,j,$ such that $\langle\psi_{i}|\phi_{j}\rangle=0$. Consider coarse-graining of all other outcomes except $i$ and $j$ for the two measurements. Then, the resulting measurements will be $\{|\psi_{i}\rangle\!\langle\psi_{i}|,\mathbbm{1}-|\psi_{i}\rangle\!\langle\psi_{i}|\}$ and $\{|\phi_{j}\rangle\!\langle\phi_{j}|,\mathbbm{1}-|\phi_{j}\rangle\!\langle\phi_{j}|\}$, which commute with each other (i.e., the resulting measurements are compatible) since $\langle\psi_{i}|\phi_{j}\rangle=0$. Thus, if the measurements are fully incompatible, condition (11) holds.
@@
III.2 Disjoint-convex-mixing of measurements
Consider a device that implements three different measurements, $M=\{M_{z}\}_{z},N=\{N_{z}\}_{z}$, and $R=\{R_{z}\}_{z}$, all having the same number of outcomes $z\in[d]$. For the first input, it performs measurement $M$ with probability $q$ and $N$ with probability $(1-q)$. For the second input, it always performs measurement $R.$ Thus, the new measurement $Q_{(M,N)}$, realized through the convex-mixing, is
$Q_{(M,N)}=\{qM_{z}+(1-q)N_{z}\}_{z},$ (14)
where $q\in[0,1]$ is the weightage of the disjoint-convex-mixing. Even if $R$ is incompatible with $M$ and $N$ separately, $R$ is not necessarily incompatible with $Q_{(M,N)}$ for all values of $q$. In a similar way, we introduce the notion of full incompatibility w.r.t. disjoint-convex-mixing.
@@ III.2 Disjoint-convex-mixing of measurements
Three measurements $M,N\,\,\text{and}\,\,R$ are fully incompatible w.r.t. disjoint-convex-mixing if each of the pairs, $M$ and $Q_{(N,R)},$ $N$ and $Q_{(M,R)},$ $R$ and $Q_{(M,N)},$ are incompatible for all values of $q$, where $Q_{(\ ,\ )}$ is defined in (14).
Consider three unbiased qubit binary outcome measurements $\{M_{0},M_{1}\},\{N_{0},N_{1}\},\{R_{0},R_{1}\}$,
$M_{z}=\frac{1}{2}\left(\mathbbm{1}+(-1)^{z}\,\,\vec{n_{0}}\cdot\vec{\sigma}\right),$
$N_{z}=\frac{1}{2}\left(\mathbbm{1}+(-1)^{z}\,\,\vec{n_{1}}\cdot\vec{\sigma}\right),$
$R_{z}=\frac{1}{2}\left(\mathbbm{1}+(-1)^{z}\,\,\vec{n_{2}}\cdot\vec{\sigma}\right),$ (15)
with $z=0,1$ and $||\vec{n_{i}}||\leqslant 1\,\,\text{where}\,\,i\in\{0,1,2\}$. A necessary and sufficient criterion for the incompatibility of two unbiased binary-outcome qubit measurements is given by Busch in Busch (1986) (also, see eq.(7) of Gühne et al. (2023)). By applying this criterion, we find that the above three measurements (III.2) are fully incompatible w.r.t. disjoint-convex-mixing if and only if,
$||\vec{n}_{i}+q\vec{n}_{j}+(1-q)\vec{n}_{k}||+||\vec{n}_{i}-q\vec{n}_{j}-(1-q)\vec{n}_{k}||>2,$ (16)
@@
Theorem 2.
If three-qubit measurements (III.2) are such that $\vec{n}_{i}$ are in the same plane of the Bloch sphere, then they are not fully incompatible w.r.t. disjoint-convex-mixing.
@@
If the three $\vec{n}_{i}$ are in the same plane, then there exists at least one triple $(i,j,k)$ such that
@@
for some $c\in(0,1],q\in[0,1]$. In other words, there exists a triple $(i,j,k)$ so that $\vec{n}_{i}$ is expressed as a linear combination of $\vec{n}_{j}$ and $\vec{n}_{k}$ with non-negative coefficients (here $q/c$ and $(1-q)/c$), where the sum of those two non-negative coefficients is greater than or equal to 1. Substituting this into left hand side of (16), we find
@@
Theorem 3.
Consider three-qubit measurements (III.2) are such that $\vec{n}_{0}=\nu_{0}\hat{x}$, $\vec{n}_{1}=\nu_{1}\hat{y}$, $\vec{n}_{2}=\nu_{2}\hat{z}$ with $0\leqslant\nu_{0},\nu_{1},\nu_{2}\leqslant 1$, that is, the noisy version of Pauli observables,
$M_{z}=\frac{1}{2}\left(\mathbbm{1}+(-1)^{z}\nu_{0}\sigma_{x}\right)=\nu_{0}\left(\frac{\mathbbm{1}+(-1)^{z}\sigma_{x}}{2}\right)+(1-\nu_{0})\frac{\mathbbm{1}}{2},$
$N_{z}=\frac{1}{2}\left(\mathbbm{1}+(-1)^{z}\nu_{1}\sigma_{y}\right)=\nu_{1}\left(\frac{\mathbbm{1}+(-1)^{z}\sigma_{y}}{2}\right)+(1-\nu_{1})\frac{\mathbbm{1}}{2},$
$R_{z}=\frac{1}{2}\left(\mathbbm{1}+(-1)^{z}\nu_{2}\sigma_{z}\right)=\nu_{2}\left(\frac{\mathbbm{1}+(-1)^{z}\sigma_{z}}{2}\right)+(1-\nu_{2})\frac{\mathbbm{1}}{2},$
with $z=0,1$. These measurements are fully incompatible w.r.t. disjoint-convex-mixing if and only if
@@
Thus, the right-land-side of (16) is greater than 2 for all values of $q$ if and only if
@@
Clearly, the three Pauli observables are fully incompatible w.r.t. disjoint-convex-mixing, and moreover, if we take an equal amount of noise $\nu_{0}=\nu_{1}=\nu_{2}=\nu$, then (20) implies $\nu>\sqrt{2/3}$.
We can generalize the notion of full incompatibility w.r.t. disjoint-convex-mixing.
Definition 4 ($k$-incompatible measurements w.r.t. disjoint-convex-mixing).
Given a set of $n$ measurements, the measurements are $k$-incompatible w.r.t. disjoint-convex-mixing if, after taking every possible disjoint-convex-mixing that yields $k$ number of measurements, the resulting measurements are incompatible.
Definition 5 (Fully incompatible measurements w.r.t. disjoint-convex-mixing).
A set of $n$ measurements is fully incompatible w.r.t. disjoint-convex-mixing if it is $k$-incompatible for all $k=2,\cdots,n$.
Observation 3.
@@
Fully incompatible measurements w.r.t. disjoint-convex-mixing imply that every pair of measurements from that set is incompatible. The reverse implication does not hold.
Proof.
@@
Consider a set of $n$ measurements, in which there is a pair of measurements $\{M_{z}\}$ and $\{N_{z}\}$ that are compatible. Now if we make two partitions where $\{M_{z}\}$ and $\{N_{z}\}$ are in different partitions and the disjoint-convex-mixing is such that the probabilities of arising all other measurements are zero, then the resultant measurement pair must be compatible. This is true for any compatible pair of measurements. Thus, if the measurements are fully incompatible w.r.t. disjoint-convex-mixing, every pair of measurements must necessarily be incompatible.
[…]
Some explicit examples of $k$-incompatible measurements w.r.t. both of these operations are provided in the subsequent section.
IV Robustness under noise for different levels of incompatibility
@@
In this section, we analyze the role of noise on quantum measurements and study how the incompatibility properties depend on it. More specifically, we study the variation of the critical amount of noise in the different layers of incompatibility under these classical operations (coarse-graining and disjoint-convex-mixing). Due to the ubiquitous nature of noise, it is pertinent to study the extent to which noise could be tolerated by a set of measurements while still retaining their incompatibility. We take a noisy version of mutually unbiased bases measurements (MUBs),
$M_{i|x}=\nu|\phi_{i|x}\rangle\langle\phi_{i|x}|+\frac{(1-\nu)}{d}\mathbbm{1}_{d},$ (23)
where $\{|\phi_{i|x}\rangle\}_{i,x}$ form mutually unbiased bases measurements in $\mathbb{C}^{d}$. Here $\nu$ is the robustness parameter (or visibility parameter) and $(1-\nu)$ is the noise parameter, $0\leq\nu\leq 1$. When the noise parameter is zero (i.e., robustness, $\nu=1$), the measurements are fully incompatible, and when the noise parameter is one (i.e., the robustness, $\nu=0$), the measurements are trivial and compatible. Our aim is to obtain the critical value of the robustness parameter above which the measurements remain incompatible before and after different classical operations.
@@
To check the compatibility, i.e., the existence of a parent POVM, we use the method described in Gühne et al. (2023). This can be cast as a semi-definite programming (SDP) problem that takes a set of measurements $\{M_{z_{x}|x}\}$ and deterministic classical post-processings $p(z_{x}|x,\lambda)$ as input, and checks whether the measurements are compatible or not, subject to the constraints
@@
where $\mu$ is the optimization parameter. This method finds the maximum value of $\mu$ for each $\{p(z_{x}|x,\lambda)\}$. If this optimization returns a negative value of $\mu$, then the constraint of Eq.(26) cannot be fulfilled, which implies that the measurements $\{M_{z_{x}|x}\}$ are incompatible. Otherwise, they are compatible.
As discussed earlier, the degree of incompatibility of measurements may be reduced due to coarse-graining of outcomes and convex mixing of measurements. In the subsequent subsection, we present an explicit analysis of how the degree of incompatibility (quantified by the robustness parameter (Heinosaari et al., 2015)) varies as the measurements are subjected to coarse-graining and disjoint-convex-mixing, in dimensions $3$ and $4$. For simplicity, we have considered the measurements to be MUBs (as discussed earlier), where the number of outcomes are same as that of their dimension. Also, as MUBs exhibit the highest degree of incompatibility, (Designolle et al., 2019b) it is a natural choice to use them for the prominent observation of the various layers of incompatibility under the classical operations. However, one can apply this formalism to any set of measurements in a similar fashion.
IV.1 Robustness of incompatibility under coarse-graining of outcomes
Dimension $3$. Let $\{M_{i}\}$ and $\{N_{j}\},i,j\in\{1,2,3\}$ be two three-outcome measurements acting on $\mathbb{C}^{3}$, where
$M_{i}=\nu|i\rangle\langle i|+(1-\nu)\frac{\mathbbm{1}}{3},$
$N_{j}=\nu|\psi_{j}\rangle\langle\psi_{j}|+(1-\nu)\frac{\mathbbm{1}}{3},$ (27)